If the RR = to 20%, and there is a new deposit of $10,000, what is...

  1. If the RR = to 20%, and there is a new deposit of $10,000, what is the maximum the money supply can expand. Show all work. 5 points

  1. If the RR = 25%, and there is a new deposit of $10,000, what is the maximum the money supply can expand. Show all work. 5 points.

  1. If the RR = 50%, and there is a new deposit of $10,000, what is the maximum the money supply can expand. Show all work. 5 points

  1. Why in questions 6 – 8 above, does the increase in the money supply decrease. 5 points

  1. If the RR = 20%, and there is a withdrawal of $20,000, what happens to the money supply? Why? Show all work. 5 points
Answer #1

Maximum increase in money supply = New deposit / RR

(1) When RR = 20% = 0.2,

Maximum increase in money supply = $10,000 / 0.2 = $50,000

(2) When RR = 25% = 0.25,

Maximum increase in money supply = $10,000 / 0.25 = $40,000

(3) When RR = 50% = 0.5,

Maximum increase in money supply = $10,000 / 0.5 = $20,000

(4) As RR increases, commercial banks are required to set aside higher proportions of new deposits as required reserves, so excess reserves available fr credit lending falls. Lower credit lending decreases money supply.

(5) A withdrawal of deposit will decrease money supply.

When RR = 20% = 0.2,

Maximum decrease in money supply = Withdrawn amount / RR = $20,000 / 0.2 = $100,000